OpenStudy (unklerhaukus):

Find eigenvectors of T \[\textbf{T} = \begin{pmatrix} 1 & 1-i \\ 1+i & 0 \\ \end{pmatrix}\] The eigenvalues are \[\lambda_{1,2}=-1,2 \in \mathbb{R}\]

OpenStudy (unklerhaukus):

\[\textbf{T}|\alpha \rangle=\lambda_1|\alpha\rangle =-1|\alpha\rangle\]

OpenStudy (unklerhaukus):

\[2\alpha_1+(1-i)\alpha_2=-\alpha_1\quad{{(i)}}\]\[(1+i)\alpha_1+\alpha_2=-\alpha_2\quad{{(ii)}}\]

OpenStudy (anonymous):

Linear algebra?

OpenStudy (unklerhaukus):

well it is , . but this question is from a Quantum Mechanics text's appendix I can find the eigenvalues, i just can't find the eigenvalues

OpenStudy (unklerhaukus):

well this is just a simple eigenvalue problem , no QM required